OpenStudy (anonymous):

Express the area and circumference of a circle as functions of the circle's radius. Then express the area as a function of the circumference.

OpenStudy (tkhunny):

You have nothing?

OpenStudy (anonymous):

I know that the formula for the area of a circle is \[A=pir ^{2}\] and the circumference is \[C=2pir\], but I don't know where to go from there.

OpenStudy (tkhunny):

Try some substitution. \(\dfrac{C}{2} = \pi r\) \(A = \pi r \cdot r\) Do you see it? It MAY mean a function of the Circumference ONLY. That's only a little more annoying. Solve the Circumference equation for 'r' and substitute the whole thing.

OpenStudy (anonymous):

so then it would be \[A=\pi r \times \frac{ C }{ 2\pi }\] which would then equal \[A=\frac{ Cr }{ 2 }\]

OpenStudy (tkhunny):

Well, maybe, but it may be more like this: \(\dfrac{C}{2} = \pi r\; and \;\dfrac{C}{2\pi} = r\) And this leads to \(A = \pi r^{2} = (\pi r)\cdot (r) = \dfrac{C}{2}\cdot\dfrac{C}{2\pi} = \dfrac{C^{2}}{4\pi} = \dfrac{1}{\pi}\cdot\left(\dfrac{C}{2}\right)^{2}\) Various ways to write it.

OpenStudy (anonymous):

alright cool. thank you!